OpenStudy (anonymous):

find the directional derivative of f at the point p in the indicated direction. f(x,y) = x^2-5xy+3y^2 p (3,-1) u= sqrt(2)/2 (i+j)

OpenStudy (anonymous):

so fx=2x-5y, fy=-5x+6y gradient of f = 11i + -21k ... i think :)

OpenStudy (anonymous):

Now, (11, -21) dot product (sqrt(2)/2, (sqrt(2)/2)

OpenStudy (anonymous):

11sqrt(2)/2 +-21sqrt(2(/2? ?

OpenStudy (anonymous):

Yes, exactly that, (a,b) dot (c,d) = ac + bd

OpenStudy (anonymous):

i got -7.06, which is pretty much the right answer.. i didn't realize it was so simple! thank you so much :D

OpenStudy (anonymous):

Besides the name, directional derivatives are easy! Good luck

OpenStudy (anonymous):

how about f(x,y)= arctan (y/x) i'm not even sure how to get the partial derivatives of that :(

OpenStudy (anonymous):

d/dx (arctan(x)) = 1/1+x^2 So...u = y/x partial x (u) = -y/x^2 d/du (arctan(u)) = 1/(1+u^2)*(-y/x^2) Replace correctly the things, CHECK my answer because im not completely sure if this makes sense, and finally repeat for partial y.

OpenStudy (anonymous):

i mean this d/du (arctan(u)) = (-y/x^2)/(1+u^2)

OpenStudy (anonymous):

forgot to replace u lol